Find the slope of the line
x-y=-2

Answer by jim_thompson5910(35256)

Solve for y

x-y=-2
x-y+y=-2+y
x=-2+y
y-2 = x
y-2+2 = x+2
y = x+2


The equation y = x+2 is in the form y = mx+b where m = 1 is the slope.

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------